OpenStudy (anonymous):

HELP HOMEWORK DUE SOON!!! Find the derivative of 4x(lnx+ln4)−6x+e

OpenStudy (anonymous):

I put 4(lnx+ln4)+4x((1/x)+(1/4))-6+e as my answer

OpenStudy (paxpolaris):

\[ 4x \left( \ln x+\ln4 \right)−6x+e\] the derivative of this whole thing...?

OpenStudy (paxpolaris):

\[4x \left( \ln x+\ln4 \right)-6x+e\]

OpenStudy (paxpolaris):

the derivative of e is 0. e is just a number

OpenStudy (paxpolaris):

same for ln(4)

OpenStudy (paxpolaris):

the derivative of ln(4) is 0 ...(not 1/4)

OpenStudy (anonymous):

thanks!

OpenStudy (paxpolaris):

so you are left with:\[4(\ln x+\ln4)+4x \left( \frac1x \right)-6\] simplify further:\[4\ln x+4\ln4-2\]